# Simple Semi-supervised Knowledge Distillation from Vision-Language Models via Dual-Head Optimization

### v1

@@ A.1 Single-Head Optimization
Theorem A.1 (Optimal Distribution for Single-Head Optimization).
@@
$-\frac{\lambda y_{c}+(1-\lambda)p_{\tau,c}}{\hat{p}_{c}}+\mu=0$ (21)
@@
A.2 Dual-Head Optimization
@@
Assumption A.2 ($\varepsilon$-Convergence).
@@
Theorem A.3 (Inference Equivalence Under $\varepsilon$-Convergence).
Under A.2, by combining the outputs of both heads as:
$\hat{p}_{\texttt{DHO}}=\alpha\cdot\sigma(h_{\text{CE}}(z))+(1-\alpha)\cdot\sigma(h_{\text{KD}}(z)/\beta),\quad\text{where }\alpha=\lambda$ (30)
@@
$\|\hat{p}_{\texttt{DHO}}-\hat{p}^{*}\|_{1}$ $=\|\alpha\cdot\sigma(h_{\text{CE}}(z))+(1-\alpha)\cdot\sigma(h_{\text{KD}}(z)/\beta)-\lambda y-(1-\lambda)p_{\tau}\|_{1}$ (32)
@@
where we applied the triangle inequality for the $\ell_{1}$ norm and used A.2.
@@
$\hat{p}_{\texttt{DHO}}\approx_{\varepsilon}\hat{p}^{*}$ (36)
@@
Lemma A.4 (Temperature Matching via KL Divergence).
@@
Corollary A.5 (Optimal DHO Configuration).
@@
$\hat{p}_{\texttt{DHO}}\approx_{\varepsilon}\hat{p}^{*}=\lambda y+(1-\lambda)p_{\tau}$ (42)
@@
Appendix B Algorithms and Implementation
B.1 Pseudocode
We present the pseudocode for DHO in Algorithms 1 and 2 for training and inference, respectively.
Algorithm 1 DHO Training with zero-shot CLIP [72] teacher
@@ B.1 Pseudocode
2:  Input: student feature extractor $g$, prediction heads $h_{\text{CE}},h_{\text{KD}}$, teacher encoders $f_{\mathbb{X}},f_{\mathbb{T}}$,
@@
13:        $p_{b}^{(l)}\leftarrow\sigma\left(\frac{1}{\zeta\cdot\eta}[\mathtt{CosSim}(f_{\mathbb{X}}(x_{b}^{(l)}),f_{\mathbb{T}}(t_{1})),\ldots,\mathtt{CosSim}(f_{\mathbb{X}}(x_{b}^{(l)}),f_{\mathbb{T}}(t_{C}))]^{\top}\right)$
@@
19:        $p_{b^{\prime}}^{(u)}\leftarrow\sigma\left(\frac{1}{\zeta\cdot\eta}[\mathtt{CosSim}(f_{\mathbb{X}}(x_{b^{\prime}}^{(u)}),f_{\mathbb{T}}(t_{1})),\ldots,\mathtt{CosSim}(f_{\mathbb{X}}(x_{b^{\prime}}^{(u)}),f_{\mathbb{T}}(t_{C}))]^{\top}\right)$
@@
2:  $z\leftarrow g(x)$
3:  $\hat{p}_{\text{CE}}\leftarrow\sigma(h_{\text{CE}}(z))$
4:  $\hat{p}_{\text{KD}}\leftarrow\sigma(h_{\text{KD}}(z)/\beta)$
5:  $\hat{p}\leftarrow\alpha\cdot\hat{p}_{\text{CE}}+(1-\alpha)\cdot\hat{p}_{\text{KD}}$
6:  $\hat{y}\leftarrow\arg\max_{c}(\hat{p}_{c})$
@@
B.2 Implementation Details
@@
Student: ResNet18 [32] from scratch or ResNet50 from DINO [8] Input size: 224$\times$224 Zero-shot Teacher: ResNet50 from CLIP [72] Few-shot Teacher: ResNet50 from Tip-Adapter-F [115] Teacher input size: 224$\times$224 labeled data: $K\in\{1,2,4,8,16\}$ shots $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: $\alpha=0.4$, $\beta=0.5$ (zero-shot); $\alpha=0.2$, $\beta=0.5$ (few-shot) Epochs: 20 Optimizer: AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) Learning rate: $1\times 10^{-3}$, weight decay: $1\times 10^{-2}$ Batch size: 512 (labeled: 256, unlabeled: 256) Scheduler: Cosine decay without warmup Augmentation: Random crops (x0.5-1.0), horizontal flips
@@
Student: ResNet18 [32] or MobileNet [77] pre-trained on ImageNet under supervision Input size: 224$\times$224 Zero-shot Teacher: ResNet50 from CLIP [72] Few-shot Teacher: ResNet50 from Tip-Adapter-F [115] Teacher input size: 224$\times$224 labeled data: $K\in\{1,2,4,8,16\}$ shots $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: determined by validation Epochs: 200 Optimizer: AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) Learning rate: $1\times 10^{-3}$, weight decay: $1\times 10^{-2}$ Batch size: 128 (labeled: 64, unlabeled: 64) Scheduler: Cosine decay without warmup Augmentation: Random crops (x0.5-1.0), horizontal flips
@@
Student: CLIP ViT-B/16 or ViT-L/14 [72] Input size: 224$\times$224 (ViT-B/16) or 336$\times$336 (ViT-L/14) Zero-shot Teacher: CLIP ViT-L/14 or ViT-H/14 [22] Teacher input size: 336$\times$336 (ViT-L/14) or 378$\times$378 (ViT-H/14) Few-shot Teacher: N/A labeled data: 1% ($\frac{N}{N+M}\approx 0.01$) or 10% ($\frac{N}{N+M}\approx 0.1$) of training data $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: $\alpha=0.5$, $\beta=0.5$ Epochs: 32 Optimizer: AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) Learning rate: $5\times 10^{-5}$, weight decay: $5\times 10^{-2}$ Batch size: 512 (labeled: 256, unlabeled: 256) Scheduler: Cosine warmup decay (5000 steps) Augmentation: Random crops (x0.5-1.0), horizontal flips
Table 6 provides a comprehensive overview of the implementation details for our three experimental settings. For ImageNet, we train ResNet50 from DINO  [8] or ResNet18 [32] from a scratch over 20 epochs with a teacher of ResNet50 from CLIP [72], while using Tip-Adapter-F [115] as few-shot teacher. In our 10 datasets experiments, ImageNet pre-trained ResNet18 [32] or MobileNet [77] serve as students with the same teachers but trained for a longer period of 200 epochs. For VLM distillation, we use smaller CLIP models (ViT-B/16, ViT-L/14) [72] as students learning from larger models (ViT-L/14 [72], ViT-H/14 [22]) over 32 epochs. All configurations use AdamW [58] optimization with appropriate learning rates, weight decay, and cosine scheduling. Table 6 also details the few-shot settings (ranging from 1-16 shots where applicable), inference parameters, and consistent data augmentation strategies across all experiments. The batch composition maintains a 1:1 ratio between labeled and unlabeled samples, i.e., $B=B^{\prime}=128$ in Algorithm 1, with $\lambda=0.5$ for loss weighting.
B.3 Computational Costs
@@
Inference overhead of DHO over SHO.
Table 7 presents computational overheads at inference time introduced by DHO over SHO for all the architectures in this paper, such as MobileNetV2 [77], ResNet-18 [32], ResNet-50 [32], ViT-B/16 [20], and ViT-L/16 [20].
@@
Table 8 presents the training time required for our experiments. For VLM distillation experiments, which represent the most resource-intensive component of our work, we used 8× NVIDIA RTX 4090 GPUs. The ViT-H/14 to ViT-L/14 distillation required approximately 80 hours, while the ViT-H/14 to ViT-B/16 and ViT-L/14 to ViT-B/16 distillations required approximately 40 and 28 hours, respectively. For the ViT-H/14 to ViT-L/14 distillation, we implemented gradient accumulation with 4 steps and mixed precision training [61] to optimize computational efficiency. For ImageNet experiments, we used 4× NVIDIA RTX 4090 GPUs, with ResNet-18 and ResNet-50 models requiring approximately 6 and 8 hours of training time, respectively. We provide these details to facilitate reproduction of our results and to give researchers a clear understanding of the computational resources needed to implement our approach at scale.
@@
Inference overhead improvements with ToMe.
To further improve the computational efficiency of our approach, we explored integrating Token Merging (ToMe) [5] with DHO. ToMe is a technique that reduces the number of tokens in ViTs by merging similar tokens to improve the efficiency of ViTs. Table 9 shows that combining DHO with ToMe significantly reduces computational costs with minimal impact on performance.
Appendix C Datasets
@@
Caltech101 [26] 100 4,128 1,649 2,465 100 (2.42%) 1,600 (38.76%)
OxfordPets [69] 37 2,944 736 3,669 37 (1.26%) 592 (20.11%)
StanfordCars [48] 196 6,509 1,635 8,041 196 (3.01%) 3,136 (48.18%)
Flowers102 [67] 102 4,093 1,633 2,463 102 (2.49%) 1,632 (39.87%)
Food101 [6] 101 50,500 20,200 30,300 101 (0.20%) 1,616 (3.20%)
FGVCAircraft [59] 100 3,334 3,333 3,333 100 (3.00%) 1,600 (48.00%)
SUN397 [99] 397 15,880 3,970 19,850 397 (2.50%) 6,352 (40.00%)
DTD [17] 47 2,820 1,128 1,692 47 (1.67%) 752 (26.67%)
EuroSAT [34] 10 13,500 5,400 8,100 10 (0.07%) 160 (1.19%)
UCF101 [81] 101 7,639 1,898 3,783 101 (1.32%) 1,616 (21.15%)
@@
ImageNet-V2 [73] 1,000 - - 10,000 - -
ImageNet-Sketch [93] 1,000 - - 50,889 - -
ImageNet-A [37] 200 - - 7,500 - -
ImageNet-R [36] 200 - - 30,000 - -
@@
We evaluated our approach on 11 diverse datasets, with ImageNet [76] serving as our primary benchmark. The datasets span general object recognition [76, 26], fine-grained classification tasks (vehicles [48, 59], natural entities [67, 69, 6]), and specialized domains (scenes [99], textures [17], remote sensing [34], and human actions [81]). Additionally, we conduct experiments on four out-of-distribution test sets to further validate the model’s generalization capabilities. To assess our model’s robustness to distribution shifts, we evaluate it on several challenging variants of ImageNet: ImageNet-v2 [73], ImageNet-Sketch [93], ImageNet-A [37], and ImageNet-R [36]. The detailed results and analyses are presented in Section D.2
@@
We summarize the overview of datasets used in Table 10, these datasets exhibit diversity in their characteristics, with varying numbers of classes and samples per dataset. This diversity enabled us to thoroughly validate our method across different few-shot semi-supervised learning scenarios by systematically varying the ratios between labeled and unlabeled samples.
Appendix D Additional Experiments
D.1 Experiments on MobileNet

### v2

@@ B.1 Single-Head Optimization
Theorem 2 (Optimal Distribution for Single-Head Optimization).
@@
$\frac{\lambda y_{c}+(1-\lambda)p_{\tau,c}}{\hat{p}_{c}}+\mu=0$ (20)
@@
B.2 Dual-Head Optimization
@@
Assumption 2 ($\varepsilon$-Convergence).
@@
Theorem 3 (Inference Equivalence Under $\varepsilon$-Convergence).
Under Assumption˜2, by combining the outputs of both heads as:
$\hat{p}_{\textbf{{DHO}}}=\alpha\cdot\sigma(h_{\text{CE}}(z))+(1-\alpha)\cdot\sigma(h_{\text{KD}}(z)/\beta),\quad\text{where }\alpha=\lambda$ (29)
@@
$\|\hat{p}_{\textbf{{DHO}}}-\hat{p}^{*}\|_{1}$ $=\|\alpha\cdot\sigma(h_{\text{CE}}(z))+(1-\alpha)\cdot\sigma(h_{\text{KD}}(z)/\beta)-\lambda y-(1-\lambda)p_{\tau}\|_{1}$ (31)
@@
where we applied the triangle inequality for the $\ell_{1}$ norm and used Assumption˜2.
@@
$\hat{p}_{\textbf{{DHO}}}\approx_{\varepsilon}\hat{p}^{*}$ (35)
@@
Lemma 1 (Temperature Matching via KL Divergence).
@@
Corollary 1 (Optimal DHO Configuration).
@@
$\hat{p}_{\textbf{{DHO}}}\approx_{\varepsilon}\hat{p}^{*}=\lambda y+(1-\lambda)p_{\tau}$ (41)
@@
Appendix C Algorithms and Implementation
C.1 Pseudocode
We present the pseudocode for DHO in Algs.˜1 and 2 for training and inference, respectively.
Algorithm 1 DHO Training with zero-shot CLIP (Radford et al., 2021) teacher
@@ C.1 Pseudocode
2:Input: student feature extractor $g$, prediction heads $h_{\text{CE}},h_{\text{KD}}$, teacher encoders $f_{\mathcal{X}},f_{\mathcal{T}}$,
@@
13:   $p_{b}^{(l)}\leftarrow\sigma\left(\frac{1}{\zeta\cdot\eta}[\mathtt{CosSim}(f_{\mathcal{X}}(x_{b}^{(l)}),f_{\mathcal{T}}(t_{1})),\ldots,\mathtt{CosSim}(f_{\mathcal{X}}(x_{b}^{(l)}),f_{\mathcal{T}}(t_{C}))]^{\top}\right)$
@@
19:   $p_{b^{\prime}}^{(u)}\leftarrow\sigma\left(\frac{1}{\zeta\cdot\eta}[\mathtt{CosSim}(f_{\mathcal{X}}(x_{b^{\prime}}^{(u)}),f_{\mathcal{T}}(t_{1})),\ldots,\mathtt{CosSim}(f_{\mathcal{X}}(x_{b^{\prime}}^{(u)}),f_{\mathcal{T}}(t_{C}))]^{\top}\right)$
@@
[…]
@@
C.2 Implementation Details
Tab.˜7 provides a comprehensive overview of the implementation details for our experiments on 1) few-shot semi-supervised settings on ImageNet and 10 datasets, 2) low-shot semi-supervised settings on ImageNet, and 3) VLM-based adaptation methods.
@@ C.2 Implementation Details
Student: ResNet18 (He et al., 2016) from scratch or ResNet50 from DINO (Caron et al., 2021) Input size: $\times224$ Zero-shot Teacher: ResNet50 from CLIP (Radford et al., 2021) Few-shot Teacher: ResNet50 from Tip-Adapter-F (Zhang et al., 2021) Teacher input size: $\times224$ labeled data: $K\in\{1,2,4,8,16\}$ shots $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: $\alpha=0.4$, $\beta=0.5$ (zero-shot); $\alpha=0.2$, $\beta=0.5$ (few-shot) Epochs: 20 Optimizer: AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) Learning rate: $1\times 10^{-3}$, weight decay: $1\times 10^{-2}$ Batch size: 512 (labeled: 256, unlabeled: 256) Scheduler: Cosine decay without warmup Augmentation: Random crops (x0.5-1.0), horizontal flips
@@
Student: ResNet18 (He et al., 2016) or MobileNet (Sandler et al., 2018) pre-trained on ImageNet under supervision Input size: $\times224$ Zero-shot Teacher: ResNet50 from CLIP (Radford et al., 2021) Few-shot Teacher: ResNet50 from Tip-Adapter-F (Zhang et al., 2021) Teacher input size: $\times224$ labeled data: $K\in\{1,2,4,8,16\}$ shots $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: determined by validation Epochs: 200 Optimizer: AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) Learning rate: $1\times 10^{-3}$, weight decay: $1\times 10^{-2}$ Batch size: 128 (labeled: 64, unlabeled: 64) Scheduler: Cosine decay without warmup Augmentation: Random crops (x0.5-1.0), horizontal flips
@@
Student: CLIP ViT-B/16 or ViT-L/14 (Radford et al., 2021) Input size: $\times224$ (ViT-B/16) or $\times336$ (ViT-L/14) Zero-shot Teacher: CLIP ViT-L/14 or ViT-H/14 (Fang et al., 2023) Teacher input size: $\times336$ (ViT-L/14) or $\times378$ (ViT-H/14) Few-shot Teacher: N/A labeled data: 1% ($\frac{N}{N+M}\approx 0.01$) or 10% ($\frac{N}{N+M}\approx 0.1$) of training data $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: $\alpha=0.5$, $\beta=0.5$ Epochs: 32 Optimizer: AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) Learning rate: $5\times 10^{-5}$, weight decay: $5\times 10^{-2}$ Batch size: 512 (labeled: 256, unlabeled: 256) Scheduler: Cosine warmup decay (5000 steps) Augmentation: Random crops (x0.5-1.0), horizontal flips
Tab.˜8 provides the implementation details for our out-of-distribution generalization experiments on 1) full training model evaluation and 2) adaptation methods including linear evaluation, visual prompt tuning, and VLM-based methods.
Full Training
Model Configuration Training Details
Student: CLIP ViT-B/16 (Radford et al., 2021) Student input size: $\times224$ Zero-shot Teacher: CLIP ViT-L/14 (Radford et al., 2021) Teacher input size: $\times336$ Labeled data: 1% and 10% ImageNet $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: $0.5$ and 11 Epochs: 32 Optimizer: AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) Learning rate: $5\times 10^{-5}$, weight decay: $5\times 10^{-2}$ Batch size: 512 (labeled: 256, unlabeled: 256) Scheduler: Cosine warmup decay (5000 steps) Augmentation: Random crops (x0.5-1.0), horizontal flips
Adaptation Methods (Linear Evaluation & Visual Prompt Tuning)
Method Configuration Training Details
Linear evaluation (Caron et al., 2021) Visual prompt tuning (Jia et al., 2022) Frozen backbone: CLIP ViT-B/16 (Radford et al., 2021) Input size: $\times224$ Zero-shot Teacher: CLIP ViT-L/14 (Radford et al., 2021) Teacher input size: $\times336$ Labeled data: 1% and 10% ImageNet $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: $0.5$ and 11 Epochs: 20 Optimizer: AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) Learning rate: $5\times 10^{-5}$, weight decay: $5\times 10^{-2}$ Batch size: 512 (labeled: 256, unlabeled: 256) Scheduler: Cosine warmup decay (5000 steps) Augmentation: Random crops (x0.5-1.0), horizontal flips
Adaptation Methods (Prompt Tuning)
Method Configuration Training Details
Prompt tuning: CoOp (Zhou et al., 2022b), PromptSRC (Khattak et al., 2023b) Frozen backbone: CLIP ViT-B/16 (Radford et al., 2021) Input size: $\times224$ Zero-shot Teacher: CLIP ViT-L/14 (Radford et al., 2021) Teacher input size: $\times336$ Labeled data: 1% and 10% ImageNet $\zeta$, $\eta$, and $\lambda$: $0.01,2,0.5$ $\alpha$ and $\beta$: $0.5$ and 11 Prompt tuning: Following PromptSRC (Khattak et al., 2023b) configurations Comparison: CasPL (Wu et al., 2025) with domain-specific unlabeled data
C.3 Computational Costs
Inference overhead of DHO over SHO.
Tab.˜9 presents computational overheads at inference time introduced by DHO over SHO for all the architectures in this paper, such as MobileNetV2 (Sandler et al., 2018), ResNet-18 (He et al., 2016), ResNet-50 (He et al., 2016), ViT-B/16 (Dosovitskiy, 2020), and ViT-L/16 (Dosovitskiy, 2020).
@@ C.3 Computational Costs
Tab.˜10 presents the training time required for our experiments. For VLM distillation experiments, which represent the most resource-intensive component of our work, we used 8× NVIDIA RTX 4090 GPUs. The ViT-H/14 to ViT-L/14 distillation required approximately 80 hours, while the ViT-H/14 to ViT-B/16 and ViT-L/14 to ViT-B/16 distillations required approximately 40 and 28 hours, respectively. For the ViT-H/14 to ViT-L/14 distillation, we implemented gradient accumulation with 4 steps and mixed precision training (Micikevicius et al., 2017) to optimize computational efficiency. For ImageNet experiments, we used 4× NVIDIA RTX 4090 GPUs, with ResNet-18 and ResNet-50 models requiring approximately 6 and 8 hours of training time, respectively. We provide these details to facilitate reproduction of our results and to give researchers a clear understanding of the computational resources needed to implement our approach at scale.
Inference overhead improvements with ToMe.
To further improve the computational efficiency of our approach, we explored integrating Token Merging (ToMe) (Bolya et al., 2022) with DHO. ToMe is a technique that reduces the number of tokens in ViTs by merging similar tokens to improve the efficiency of ViTs. Tab.˜11 shows that combining DHO with ToMe significantly reduces computational costs with minimal impact on performance.
@@
Appendix D Datasets
@@
Caltech101 (Fei-Fei et al., 2004) 100 4,128 1,649 2,465 100 (2.42%) 1,600 (38.76%)
OxfordPets (Parkhi et al., 2012) 37 2,944 736 3,669 37 (1.26%) 592 (20.11%)
StanfordCars (Krause et al., 2013) 196 6,509 1,635 8,041 196 (3.01%) 3,136 (48.18%)
Flowers102 (Nilsback & Zisserman, 2008) 102 4,093 1,633 2,463 102 (2.49%) 1,632 (39.87%)
Food101 (Bossard et al., 2014) 101 50,500 20,200 30,300 101 (0.20%) 1,616 (3.20%)
FGVCAircraft (Maji et al., 2013) 100 3,334 3,333 3,333 100 (3.00%) 1,600 (48.00%)
SUN397 (Xiao et al., 2010) 397 15,880 3,970 19,850 397 (2.50%) 6,352 (40.00%)
DTD (Cimpoi et al., 2014) 47 2,820 1,128 1,692 47 (1.67%) 752 (26.67%)
EuroSAT (Helber et al., 2019) 10 13,500 5,400 8,100 10 (0.07%) 160 (1.19%)
UCF101 (Soomro, 2012) 101 7,639 1,898 3,783 101 (1.32%) 1,616 (21.15%)
@@
ImageNet-V2 (Recht et al., 2019) 1,000 - - 10,000 - -
ImageNet-Sketch (Wang et al., 2019) 1,000 - - 50,889 - -
ImageNet-A (Hendrycks et al., 2021b) 200 - - 7,500 - -
ImageNet-R (Hendrycks et al., 2021a) 200 - - 30,000 - -
@@
We evaluated our approach on 11 diverse datasets, with ImageNet (Russakovsky et al., 2015) serving as our primary benchmark. The datasets span general object recognition (Russakovsky et al., 2015; Fei-Fei et al., 2004), fine-grained classification tasks (vehicles (Krause et al., 2013; Maji et al., 2013), natural entities (Nilsback & Zisserman, 2008; Parkhi et al., 2012; Bossard et al., 2014)), and specialized domains (scenes (Xiao et al., 2010), textures (Cimpoi et al., 2014), remote sensing (Helber et al., 2019), and human actions (Soomro, 2012)). Additionally, we conduct experiments on four out-of-distribution test sets to further validate the model’s generalization capabilities. To assess our model’s robustness to distribution shifts, we evaluate it on several challenging variants of ImageNet: ImageNet-v2 (Recht et al., 2019), ImageNet-Sketch (Wang et al., 2019), ImageNet-A (Hendrycks et al., 2021b), and ImageNet-R (Hendrycks et al., 2021a).
@@
We summarize the overview of datasets used in Tab.˜12, these datasets exhibit diversity in their characteristics, with varying numbers of classes and samples per dataset. This diversity enabled us to thoroughly validate our method across different few-shot semi-supervised learning scenarios by systematically varying the ratios between labeled and unlabeled samples.
Appendix E Additional Experiments
E.1 Experiments on MobileNet